# I-SHEEP: Self-Alignment of LLM from Scratch through an Iterative Self-Enhancement Paradigm

### v2

@@ Abstract
Large Language Models (LLMs) have achieved significant advancements, however, the common learning paradigm treats LLMs as passive information repositories, neglecting their potential for active learning and alignment. Some approaches train LLMs using their own generated synthetic data, exploring the possibility of active alignment. However, there is still a huge gap between these one-time alignment methods and the continuous automatic alignment of humans. In this paper, we introduce I-SHEEP, an Iterative Self-EnHancEmEnt Paradigm. This human-like paradigm enables LLMs to continuously self-align from scratch with nothing. Compared to the one-time alignment method Dromedary (Sun et al. 2023b), which refers to the first iteration in this paper, I-SHEEP can significantly enhance capacities on both Qwen and Llama models. I-SHEEP achieves a maximum relative improvement of 78.2% in the Alpaca Eval, 24.0% in the MT Bench, and an absolute increase of 8.88% in the IFEval accuracy over subsequent iterations in Qwen-1.5 72B model. Additionally, I-SHEEP surpasses the base model in various standard benchmark generation tasks, achieving an average improvement of 24.77% in code generation tasks, 12.04% in TrivialQA, and 20.29% in SQuAD. We also provide new insights based on the experiment results. Our codes, datasets, and models are available at https://github.com/multimodal-art-projection/I-SHEEP.
Introduction
Large Language Models (LLMs) have achieved significant success, yet they remain far from achieving the autonomous self-regulation and coherent self-understanding characteristic of human-like Artificial General Intelligence (AGI). In the pretraining phase, LLMs learn and memorize common knowledge from vast amounts of raw text. During the Supervised Fine-Tuning (SFT) phase, LLMs are trained with question-answer pairs to develop instruction-following capabilities. Both phases currently treat LLMs as passive information repositories, neglecting their potential for active learning and alignment. Furthermore, aligning a base model from scratch requires substantial high-quality data across various tasks and scenarios. This data collection, annotation, and cleaning process is labor-intensive, time-consuming, and must ensure diversity and representativeness. Humans can automatically answer relevant questions after acquiring knowledge (Khader et al. 2016), and there is no need for an explicit process like SFT to teach humans how to answer questions. Therefore, exploring active, automatic, and continuous alignment from scratch, akin to human learning, is a crucial step for LLMs toward achieving AGI.
There are several approaches developed to align base models in low-resource scenarios, aiming to reduce or eliminate reliance on human supervision signals (Wang et al. 2022b; Sun et al. 2023b, a; Xu et al. 2024b). For instance, Self Instruct and Dromedary (Wang et al. 2022b; Sun et al. 2023b) actively improve themselves using self-generated data, while Magpie (Xu et al. 2024b) aligns the base model by prompting more powerful models to generate instruction-output pair data. Although these methods demonstrate some level of proactivity, there remains a significant gap between these one-time alignment processes and the continuous, automatic alignment for various environments seen in human learning.
@@ Introduction
Educational research suggests that metacognitive self-assessment plays a vital role in continuous alignment, helping students reflect on their knowledge and skills, manage cognitive resources, and improve their performance (Yan et al. 2023). Inspired by this perspective, we explore a new human-like paradigm for LLMs, I-SHEEP, designed to enable LLMs to proactively, automatically, and continuously align from scratch, even with minimal external signal. As shown in Figure 1, I-SHEEP begins with seed data and leverages LLMs powerful understanding and generation capabilities to create additional instruction-output pair data. We then perform self-assessment, allowing the model to monitor and assess its learning process. By filtering out incorrect cognitions and retaining accurate ones, LLMs can self-align by training themselves with these correct cognitions. Through iterative repetition of this process, the model can continuously and autonomously align from scratch, relying solely on its internal knowledge.
The main contributions can be summarized as follows: (1) We propose I-SHEEP, a human-like learning paradigm for LLMs, enabling active, automatic, and continuous self-alignment from scratch using only their internal knowledge. Based on the one-time alignment approach, I-SHEEP explores the incorporation of metacognitive self-assessment to monitor and manage the learning process as well as continuous self-alignment from scratch. (2) We integrate metacognitive self-assessment into the alignment process, allowing the model to monitor and manage its learning process. Experiments explore how varying levels of metacognitive abilities impact the efficiency of the I-SHEEP framework. (3) We find that the self-enhancement potential is closely related to the metacognitive self-assessment level and model size. We conducted experiments to explore the performance of I-SHEEP on different model sizes and varying levels of metacognitive self-assessment.
Related Work
Automatic Data Selection
@@
Zhou et al.; Bai et al. emphasize that dataset quality outweighs quantity during the instruction fine-tuning stage. As a result, some studies on instruction data selection have emerged, focusing on identifying high-quality subsets from candidate datasets(Li et al. 2023a; Du, Zong, and Zhang 2023; Liu et al. 2023; Li et al. 2024; Ge et al. 2024; Xia et al. 2024). These methods aim to improve the model performance, accelerate the training process, and facilitate data-efficient alignment. Li et al. introduce an Instruction-Following Difficulty (IFD) metric and use it to select the top 5% of data for fine-tuning models. The filtering phase in the I-SHEEP framework does not rely on carefully selected metrics, external models or human assistance. Additionally, our method is orthogonal to these selection approaches.
Synthetic Data for Improving Model
@@
Generating synthetic data refers to using the powerful generative capabilities of LLMs to create new data that simulates potential real-world scenarios, reducing the need for costly manual labeling. Some methods use the model’s self-generated data to improve itself (Wang et al. 2022b; Sun et al. 2023b, a; Yehudai et al. 2024). Other methods leverage powerful closed models to generate synthetic data, enhancing the capabilities of open-source models (Taori et al. 2023; Chiang et al. 2023; Xu et al. 2023a; Yu et al. 2023; Wei et al. 2023). In addition to generating complete instruction-output pairs, some methods collect existing raw data and synthesize corresponding questions or answers to create supervised data for improving the model (Huang et al. 2022; Li et al. 2023b; Zheng et al. 2024b; Mitra et al. 2024; Wang et al. 2022a; Asai et al. 2023). Some methods begin with instruction-output pairs, generating feedback or refining answers to improve data quality and enhance the model’s reasoning capabilities.(Lu et al. 2023; Li and He 2024; Gou et al. 2023). The I-SHEEP framework evolves from the aforementioned static, one-time improvement paradigm to a dynamic, continuous self-enhancement process.
Iterative Enhancement for LLMs
@@
There are several approaches to iterative enhancement that rely on the help of strong models or external tools (Chen et al. 2024, 2023; Lu et al. 2023; Gao et al. 2023; Lee et al. 2024). IterAlign (Chen et al. 2024) employs strong models like GPT-4 and Claude2 to detect and correct errors in responses from base LLMs and give the corresponding constitution for improving the safety of LLMs. These methods in iterative enhancement typically depend on strong models or external tools to guarantee ongoing model optimization and avoid model collapse. In addition, some methods explore iterative enhancement in the RLHF phase to continuously align the model with human preference (Yuan et al. 2024; Liu et al. 2024; Pang et al. 2024; Xu et al. 2024a, 2023b; Wu et al. 2024; Wang et al. 2024). These iterative RLHF methods start with the aligned model, while we focus on the base model continuous self-alignment from scratch.
Methodology
Self-Driven Data Synthesis
@@
Self Instruct (Wang et al. 2022b) leverages an off-the-shelf large language model (LLM) for the generation of synthetic data. The approach starts with a small set of 175 prompts, known as the seed task pool, leveraging the model’s powerful understanding and generative capabilities to generate a broader range of prompts and responses. This section elaborates on the Self-Driven Data Synthesis process from two perspectives: Instruction generation and response generation. For ease and consistency in data creation, we utilize a standardized instruction format introduced by Alpaca (Taori et al. 2023), enabling the direct generation of instructions along with their corresponding potential inputs.
@@
Instruction generation. Having some prompts from the seed dataset and the meta-prompt from Alpaca (Taori et al. 2023). The process that model generating new prompt set through In-Context Learning (ICL) can be modeled as:
@@
Self-Assessment and Data Filtering
@@
Data Filtering. After the self-assessment, the subsequent data filtering phase discards entries that do not meet the specified quality threshold. This step guarantees that only entries of the highest quality are retained in the dataset, thereby enhancing the overall reliability and utility of the generated data. Initially, we apply heuristic rule-based filtering to the generated data during data generation, following the Self-Instruct (Wang et al. 2022b). Additionally, after data generation, we filter the instruction-output pairs based on the assessment scores from the self-assessment phrase. A threshold is applied to filter based on assessment scores, yielding a high-quality dataset .
@@
Iterative Continuous Model Enhancements
Experiments
Evaluation
Chat Evaluation
We evaluate the instruction-following ability and response quality of aligned models with three chat benchmarks, AlpacaEval(Dubois et al. 2023), MT-Bench(Zheng et al. 2024a), and IFEval(Zhou et al. 2023), due to their comprehensiveness, fine granularity, and reproducibility. Both AlpacaEval and MT-Bench rely on GPT as an evaluator. IFEval provides four types of accuracy scores: prompt-level strict-accuracy, inst-level strict-accuracy, prompt-level loose-accuracy, and inst-level loose-accuracy.
OpenCompass Evaluation
We use the OpenCompass evaluation platform (Contributors 2023), a comprehensive one-stop platform for LLM evaluation. The evaluation includes standard benchmarks such as BoolQ (Clark et al. 2019), PIQA (Bisk et al. 2019), SIQA (Sap et al. 2019), HellaSwag (Zellers et al. 2019), WinoGrande (Sakaguchi et al. 2019), ARC-c (Clark et al. 2018), OpenBookQA-Fact (Mihaylov et al. 2018), CommonsenseQA (Contributors 2023), and MMLU (Hendrycks et al. 2020). It also includes code generation benchmarks such as HumanEval (Chen et al. 2021) and MBPP (Austin et al. 2021), word knowledge benchmark TriviaQA (Joshi et al. 2017), and reading comprehension benchmark SQuAD2.0 (Rajpurkar, Jia, and Liang 2018). Full results on these benchmarks are available in Appendix CC.
Main Settings
We conduct experiments on the Qwen-1.5 (Team 2024) and Llama-3 (Dubey et al. 2024) models to validate the effectiveness and generalization of I-SHEEP. Additionally, we explore the impact of different model sizes on I-SHEEP by conducting experiments on Qwen-1.5 1.8B, 4B, 7B, 14B, 32B, and 72B models, providing a detailed analysis based on the experimental results. In each iteration, the dataset for training is generated by the model from the last iteration. The case study of the generated data and the overall quality analysis can be found in Appendix BB and Appendix FF, respectively. We utilized LLaMA-Factory (Zheng et al. 2024c) for LoRA fine-tuning, with specific parameters detailed in Appendix EE. Under the configuration of using VLLM for inference (Kwon et al. 2023), the maximum duration of each iteration is about 4 hours on NVIDIA A800-SXM4-80GB8, equivalent to one iteration time for Qwen-1.5 72B.
Self-Assessment and Filter Settings
During the self-assessment phase, we propose three variants, simple standard prompt, combined standard prompt, and ICL prompt, to evaluate data quality. Detailed prompt contents can be found in appendix AA.
@@
Baseline
We use the base model, Self Instruct (Wang et al. 2022b), and Dromedary (Sun et al. 2023b) as baselines to explore the continuous and automatic enhancement of the human-like framework, I-SHEEP. Self Instruct is a one-time alignment approach where LLMs are trained directly on data they generate, without a self-assessment phase. Similarly, Dromedary is a one-time alignment process where the model generates responses following specific principles, which are then engraved into the model. This approach is similar to the first iteration setting described in this paper.
Iterative Settings and Ablation Settings
Iterative Settings. We investigate the impact of I-SHEEP on efficiency across different iterative self-enhancement settings, including using data generated by the last iteration model to train the base model, using data generated by the last iteration model to train the last iteration model, and using data generated by all previous iterations to train the base model. Additionally, we directly generate 20K and 30K data points for comparative experiments to eliminate the influence of data size in the iterative settings mentioned above. Notably, in the first iteration, all settings are identical, where the base model generates 10k data, filters it, and uses it to fine-tune itself, akin to the Dromedary(Sun et al. 2023b).
@@
Results
Main Results
Table 1 shows the experimental performance of various model sizes across different iteration steps. There are some new findings: (1) I-SHEEP exhibits efficacy across various model sizes, with particularly notable improvements in 72B. I-SHEEP achieves a maximum relative improvement of 78.2% in the Alpaca Eval, 24.0% in the MT Bench, and an absolute increase of 8.88% in the IFEval prompt-level strict accuracy over subsequent iterations in Qwen-1.5 72B model. Additionally, I-SHEEP surpasses the base model in various standard benchmark generation tasks, achieving an average improvement of 24.77% in code generation tasks, 12.04% in Trivial QA, and 20.29% in SQuAD. we find that the scores for the second round of dialogues drop significantly after the fourth iteration. This decline is likely due to our generated data consisting solely of single-round dialogues, which do not improve and may even harm the scores for the second round of dialogues. More analysis can be found in the Appendix DD. (2) The potential for improvement varies with different model sizes. The 1.8B, 4B, 7B, and 14B models exhibit improvements over two iterations, 32B and 72B model can improve three and five iterations, respectively, according to the IFEval benchmark.
Iterative Setting Results

### v3

@@ Abstract
Large Language Models (LLMs) have achieved significant advancements, however, the common learning paradigm treats LLMs as passive information repositories, neglecting their potential for active learning and alignment. Some approaches train LLMs using their own generated synthetic data, exploring the possibility of active alignment. However, there is still a huge gap between these one-time alignment methods and the continuous automatic alignment of humans. In this paper, we introduce I-SHEEP, an Iterative Self-EnHancEmEnt Paradigm. This human-like paradigm enables LLMs to iteratively self-improve even in low-resource scenarios. Compared to the one-time alignment method Dromedary Sun et al. (2023b), which refers to the first iteration in this paper, I-SHEEP can significantly enhance capacities on both Qwen and Llama models. I-SHEEP achieves a maximum relative improvement of 78.2% in the Alpaca Eval, 24.0% in the MT Bench, and an absolute increase of 8.88% in the IFEval accuracy over subsequent iterations in Qwen-1.5 72B model. Additionally, I-SHEEP surpasses the base model in various standard benchmark generation tasks, achieving an average improvement of 24.77% in code generation tasks, 12.04% in TrivialQA, and 20.29% in SQuAD. We also provide new insights based on the experiment results. Our code, datasets, and models are available at https://anonymous.4open.science/r/SHEEP/.
Yiming Liang1,2,8††thanks: Equal contribution. Ge Zhang3,5,611footnotemark: 1 Xingwei Qu4,5,611footnotemark: 1 Tianyu Zheng5,611footnotemark: 1 Jiawei Guo5,6 Xinrun Du5,6 Zhenzhu Yang5 Jiaheng Liu5 Chenghua Lin4 Lei Ma7,8 Wenhao Huang6††thanks: Corresponding authors. Jiajun Zhang1,222footnotemark: 2 1School of Artificial Intelligence, University of Chinese Academy of Sciences 2Institute of Automation, Chinese Academy of Sciences, 3University of Waterloo 4The University of Manchester, 5M-A-P, 601.ai, 7Peking University, 8BAAI
1 Introduction
Early studies improve model performance using human-labeled data, but the high cost of labeling limits scalability Zhou et al. (2024); Zheng et al. (2024b). Some methods use powerful models to synthesize data, thereby improving student models Taori et al. (2023); Xu et al. (2024b). However, these methods face performance ceilings and indirectly depend on strong models’ reliance on human-labeled signals Li et al. (2023b). Additionally, they often treat models as passive information repositories, overlooking the models’ ability to actively align. Other methods focus on the active alignment capabilities of LLMs, enhancing them through self-generated data. Nevertheless, these approaches typically rely on substantial external signals or tools, such as raw text Li et al. (2023b), retrieval-augmented generation (RAG) Asai et al. (2023), feedback from strong models Lee et al. (2024), and high-quality questions Huang et al. (2022); Wang et al. (2024), to achieve self-improvement.
Recently, some approaches explore the active alignment capabilities of LLMs in low-resource scenarios, aiming for models to self-improve with minimal reliance on external signals Wang et al. (2022b); Sun et al. (2023b, a). For example, Self-Instruct Wang et al. (2022b) prompts the model to generate instructions using a seed dataset containing only 175 human-labeled instruction pairs, achieving self-alignment. Dromedary Sun et al. (2023b) uses 16 manually crafted principles to guide LLMs in generating instruction pair data, enhancing the quality of synthesized data. However, these methods are typically one-time alignment approaches, showing significant gaps compared to the continuous and automatic alignment that humans perform in varying environments. In this paper, we explore leveraging the model’s internal metacognitive self-assessment to enable multi-round iterative self-improvement in low-resource settings, similar to human processes.
@@ 1 Introduction
Educational research suggests that metacognitive self-assessment plays a vital role in continuous alignment, helping students reflect on their knowledge and skills, manage cognitive resources, and improve their performance Yan et al. (2023). Inspired by this perspective, we introduce I-SHEEP, a human-like paradigm that enables LLMs to iteratively self-improve in low-resource settings. As shown in Figure 1, I-SHEEP begins with seed data and leverages the understanding and generation capabilities of LLMs to create additional instruction pairs. We then perform self-assessment, allowing LLMs to monitor and assess their learning process. By filtering out incorrect cognitions and retaining accurate ones, LLMs can self-improve by aligning themselves with these correct cognitions. Through an iterative process, LLMs can continuously self-align, relying solely on their internal knowledge.
The main contributions can be summarized as follows: (1) We introduce I-SHEEP, which aims to explore the potential of LLMs to iteratively self-improve in low-resource scenarios. I-SHEEP incorporates metacognitive self-assessment to monitor and manage the learning process of LLMs, enabling iterative self-improvement. (2) We analyze the factors that influence the continuous improvement potential of LLMs. Our experiments show that the self-improvement ability of LLMs is influenced by their inherent capabilities and metacognitive levels, and varies with model size and metacognitive capacity. (3) We validate the effectiveness and efficiency of the I-SHEEP framework through experiments. Even in low-resource scenarios, I-SHEEP significantly improves the performance of LLMs on various chat benchmarks and standard benchmarks through multiple iterations.
2 Related Work
2.1 Automatic Data Selection
@@
Zhou et al.; Bai et al. emphasize that dataset quality outweighs quantity during the instruction fine-tuning stage. As a result, some studies on instruction data selection have emerged, focusing on identifying high-quality subsets from candidate datasetsLi et al. (2023a); Du et al. (2023); Liu et al. (2023); Li et al. (2024); Ge et al. (2024); Xia et al. (2024). These methods aim to improve the model performance, accelerate the training process, and facilitate data-efficient alignment. Li et al. introduce an Instruction-Following Difficulty (IFD) metric and use it to select the top 5% of data for fine-tuning models. The filtering phase in the I-SHEEP framework does not rely on predefined metrics, external models, or human assistance, and is orthogonal to existing selection methods.
2.2 Synthetic Data for Improving Model
@@
Generating synthetic data refers to using the powerful generative capabilities of LLMs to create new data that simulates potential real-world scenarios, reducing the need for costly manual labeling. Some methods use the model’s self-generated data to improve itself Wang et al. (2022b); Sun et al. (2023b, a); Yehudai et al. (2024). Other methods leverage powerful closed models to generate synthetic data, enhancing the capabilities of open-source models Taori et al. (2023); Chiang et al. (2023); Xu et al. (2023a); Yu et al. (2023); Wei et al. (2023). In addition to generating complete instruction-output pairs, some methods collect existing raw data and synthesize corresponding questions or answers to create supervised data for improving the model Huang et al. (2022); Li et al. (2023b); Zheng et al. (2024b); Mitra et al. (2024); Wang et al. (2022a); Asai et al. (2023). Some methods begin with instruction-output pairs, generating feedback or refining answers to improve data quality and enhance the model’s reasoning capabilities.Lu et al. (2023); Li and He (2024); Gou et al. (2023). The I-SHEEP framework evolves from the aforementioned static, one-time improvement paradigm to a dynamic, continuous self-enhancement process.
2.3 Iterative Enhancement for LLMs
@@
There are several approaches to iterative enhancement that rely on the help of strong models or external tools Chen et al. (2024, 2023); Lu et al. (2023); Gao et al. (2023); Lee et al. (2024). IterAlign Chen et al. (2024) employs strong models like GPT-4 and Claude2 to detect and correct errors in responses from base LLMs and give the corresponding constitution for improving the safety of LLMs. These methods in iterative enhancement typically depend on strong models or external tools to guarantee ongoing model optimization and avoid model collapse. In addition, some methods explore iterative enhancement in the RLHF phase to continuously align the model with human preference Yuan et al. (2024); Liu et al. (2024); Pang et al. (2024); Xu et al. (2024a, 2023b); Wu et al. (2024); Wang et al. (2024). These iterative RLHF methods start with the aligned model, while we focus on the base model continuous self-alignment from scratch.
3 Methodology
3.1 Self-Driven Data Synthesis
@@
Self Instruct Wang et al. (2022b) leverages an off-the-shelf large language model (LLM) for the generation of synthetic data. The approach starts with a small set of 175 prompts, known as the seed task pool, leveraging the model’s powerful understanding and generative capabilities to generate a broader range of prompts and responses. This section elaborates on the Self-Driven Data Synthesis process from two perspectives: Instruction generation and response generation. For ease and consistency in data creation, we utilize a standardized instruction format introduced by Alpaca Taori et al. (2023), enabling the direct generation of instructions along with their corresponding potential inputs.
@@
Instruction generation. Having some prompts from the seed dataset and the meta-prompt from Alpaca Taori et al. (2023). The process that model generating new prompt set through In-Context Learning (ICL) can be modeled as:
@@
3.2 Self-Assessment and Data Filtering
@@
Data Filtering. After the self-assessment, the subsequent data filtering phase discards entries that do not meet the specified quality threshold. This step guarantees that only entries of the highest quality are retained in the dataset, thereby enhancing the overall reliability and utility of the generated data. Initially, we apply heuristic rule-based filtering to the generated data during data generation, following the Self-Instruct Wang et al. (2022b). Additionally, after data generation, we filter the instruction-output pairs based on the assessment scores from the self-assessment phrase. A threshold is applied to filter based on assessment scores, yielding a high-quality dataset .
3.3 Iterative Model Enhancements
4 Experiments
4.1 Evaluation
4.1.1 Chat Evaluation
We evaluate the instruction-following ability and response quality of aligned models with three chat benchmarks, AlpacaEvalDubois et al. (2023), MT-BenchZheng et al. (2024a), and IFEvalZhou et al. (2023), due to their comprehensiveness, fine granularity, and reproducibility. Both AlpacaEval and MT-Bench rely on GPT as an evaluator. IFEval provides four types of accuracy scores: prompt-level strict-accuracy, inst-level strict-accuracy, prompt-level loose-accuracy, and inst-level loose-accuracy.
@@ 4.1.1 Chat Evaluation
4.1.2 OpenCompass Evaluation
We use the OpenCompass evaluation platform Contributors (2023), a comprehensive one-stop platform for LLM evaluation. The evaluation includes standard benchmarks such as BoolQ Clark et al. (2019), PIQA Bisk et al. (2019), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2019), ARC-c Clark et al. (2018), OpenBookQA-Fact Mihaylov et al. (2018), CommonsenseQA Contributors (2023), and MMLU Hendrycks et al. (2020). It also includes code generation benchmarks such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021), word knowledge benchmark TriviaQA Joshi et al. (2017), and reading comprehension benchmark SQuAD2.0 Rajpurkar et al. (2018). Full results on these benchmarks are available in Appendix C.
4.2 Main Settings
We conduct experiments on the Qwen-1.5 Team (2024) and Llama-3 Dubey et al. (2024) models to validate the effectiveness and generalization of I-SHEEP. Additionally, we explore the impact of different model sizes on I-SHEEP by conducting experiments on Qwen-1.5 1.8B, 4B, 7B, 14B, 32B, and 72B models, providing a detailed analysis based on the experimental results. In each iteration, the dataset for training is generated by the model from the last iteration. The case study of the generated data and the overall quality analysis can be found in Appendix B and Appendix F, respectively. We utilized LLaMA-Factory Zheng et al. (2024c) for LoRA fine-tuning, with specific parameters detailed in Appendix E. Under the configuration of using VLLM for inference Kwon et al. (2023), the maximum duration of each iteration is about 4 hours on NVIDIA A800-SXM4-80GB8, equivalent to one iteration time for Qwen-1.5 72B.
4.3 Self-Assessment and Filter Settings
During the self-assessment phase, we propose three variants, simple standard prompt, combined standard prompt, and ICL prompt, to evaluate data quality. Detailed prompt contents can be found in Appendix A.
@@ 4.3 Self-Assessment and Filter Settings
4.4 Baseline
We use the base model, Self Instruct Wang et al. (2022b), and Dromedary Sun et al. (2023b) as baselines to explore the continuous and automatic enhancement of the human-like framework, I-SHEEP. Self Instruct is a one-time alignment approach where LLMs are trained directly on data they generate, without a self-assessment phase. Similarly, Dromedary is a one-time alignment process where the model generates responses following specific principles, which are then engraved into the model. This approach is similar to the first iteration setting described in this paper.
4.5 Iterative Settings and Ablation Settings
Iterative Settings. We investigate the impact of I-SHEEP on efficiency across different iterative self-enhancement settings, including using data generated by the last iteration model to train the base model, using data generated by the last iteration model to train the last iteration model, and using data generated by all previous iterations to train the base model. Additionally, we directly generate 20K and 30K data points for comparative experiments to eliminate the influence of data size in the iterative settings mentioned above. Notably, in the first iteration, all settings are identical, where the base model generates 10k data, filters it, and uses it to fine-tune itself, akin to the DromedarySun et al. (2023b).
@@ 4.5 Iterative Settings and Ablation Settings
5 Results
5.1 Main Results
Table 1 shows the experimental performance of various model sizes across different iteration steps. There are some new findings: (1) I-SHEEP exhibits efficacy across various model sizes, with particularly notable improvements in 72B. I-SHEEP achieves a maximum relative improvement of 78.2% in the Alpaca Eval, 24.0% in the MT Bench, and an absolute increase of 8.88% in the IFEval prompt-level strict accuracy over subsequent iterations in Qwen-1.5 72B model. Additionally, I-SHEEP surpasses the base model in various standard benchmark generation tasks, achieving an average improvement of 24.77% in code generation tasks, 12.04% in Trivial QA, and 20.29% in SQuAD. we find that the scores for the second round of dialogues drop significantly after the fourth iteration. This decline is likely due to our generated data consisting solely of single-round dialogues, which do not improve and may even harm the scores for the second round of dialogues. More analysis can be found in the Appendix D. (2) The potential for improvement varies with different model sizes. The 1.8B, 4B, 7B, and 14B models exhibit improvements over two iterations, 32B and 72B model can improve three and five iterations, respectively, according to the IFEval benchmark.
5.2 Iterative Setting Results